OpenStudy (nuccioreggie):

What is the value of x? A right angle is shown divided into two parts. The measure of one part of the right angle is 10 degrees. The measure of the other part is 2x 20 40 45 85

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

we can write: 2x +10 = 90 what is x?

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

hint: 2 x= 90-10= 80

OpenStudy (nuccioreggie):

Its D?

OpenStudy (nuccioreggie):

??

OpenStudy (nuccioreggie):

Is it D is D. The correct answer?

OpenStudy (michele_laino):

no, what is: x=80/2=...?

OpenStudy (nuccioreggie):

Ok so its B 40

OpenStudy (michele_laino):

that's right!